OpenStudy (s_emily):

HELP?! I have a HUGE test on Friday and this is a practice question. Walk me through it?! A display case has room for 10 statues. Find the number of ways 10 statues can be displayed if each number of statues is available. 1.) (A.) 10 statues (B.) 20 statues

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if i read the question correctly it is \(10!\) a rather large number

OpenStudy (s_emily):

my teachers had a meeting so no one could help since all the subs know nothing about Geometry. Advanced class

OpenStudy (misty1212):

actually this has nothing to do with geometry it is "combinations and permutations"

OpenStudy (s_emily):

i got 10! for A But don't know what to do for B

OpenStudy (misty1212):

i don't understand B does B mean "20 statues" instead of ten?`

OpenStudy (misty1212):

ooh actually i do sorry

OpenStudy (misty1212):

if there are 20 statues, and only ten spots to put them in, then \[20\times 19\times 18\times 17\times 16\times 15\times 14\times 13\times 12\times 11\]

OpenStudy (s_emily):

so 20 factorial?

OpenStudy (misty1212):

no

OpenStudy (misty1212):

\(20!\) has more numbers lets go slow

OpenStudy (s_emily):

oh i see only 10 slots 20x19x18x17x16x15x14x13x12x11 = ____

OpenStudy (misty1212):

you have 20 statues but only oh ok good

OpenStudy (s_emily):

ty

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]